OpenStudy (anonymous):

1. What is the colour of FeSO4.7H20 crystals?How does this colour change upon heating? Give balanced chemical equation for the changes. 2. a. What are redox reactions? b.Why is the reaction between MnO2 and HCl a redox reaction? c.Identify the substance oxidized and the substance reduced in the above reaction. 3. Balance the chemical equation:- Fe(s) + H2O(g)=Fe3O4(s) + H2(g) 4. Explain why: a)Respiration is an exothermic reaction. b)All decomposition reactions are endothermic reactions.

sam (.sam.):

Melanterite or FeSO4.7H20 is green in color

sam (.sam.):

2FeSO4 ---> Fe2O3 + SO2 + SO3

OpenStudy (anonymous):

thank you for your guidance .Sam.

sam (.sam.):

Redox reactions include all chemical reactions in which atoms have their oxidation state changed

sam (.sam.):

2b, \[\begin{array}{c} \text{Mn}\text{O}_2 \\\end{array}\text{ + }\begin{array}{cc} 4 & \text{H}\text{Cl} \\\end{array}\text{ $\longrightarrow $ }\begin{array}{c} \text{Mn}\text{Cl}_2 \\\end{array}\text{ + }\begin{array}{cc} 2 & \text{H}_2\text{O} \\\end{array}\text{ + }\begin{array}{c} \text{Cl}_2 \\\end{array}\] You can see that Mn in \(\text{Mn}\text{Cl}_2\) charge is +2, But in MnO2, Mn is +4, so, Mn charge is reduced from +4 to +2

sam (.sam.):

Did you learn Disproportionation?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Thank you for the guidance.

sam (.sam.):

Welcome. The Cl should be oxidized . For 3, 3 Fe + 4 H2O = Fe3O4 + 4 H2

OpenStudy (anonymous):

Can you please tell me in question 1 , what is the colour of the crystals after heating?Also why does it change?

OpenStudy (anonymous):

I am so sorry , it is 'how' does it change upon heating not 'why'.

sam (.sam.):

If you heat it , it will lose water and became FeSO4

OpenStudy (anonymous):

and what will be the colour after heating?